# Learning Complete Topology-Aware Correlations Between Relations for Inductive Link Prediction

### v2

@@ 3.5.3 The CCN subgraph method
[…]
@@ 3.5.4 The CCN+ subgraph method
[…]
@@ 4.4.2 Ranking metric
TABLE VI: Ablation results of triples classification on the inductive benchmark datasets extracted from WN18RR, FB15k-237, and NELL-995. “${}_{+}$ w/o RA” represent the baseline that omits the relation aggregation in ${}_{+}$. “${}_{+}$ w/o RC” represent the baseline that performs relation aggregation without modeling correlations between relations in ${}_{+}$.
@@
${}_{+}$ w/o RA 97.75 97.33 91.95 97.64 92.06 95.83 97.05 97.42 98.06 97.33 96.46 98.26
${}_{+}$ w/o RC 98.82 93.17 90.29 97.85 92.39 95.97 96.69 99.22 98.73 97.32 98.81 98.39
${}_{+}$ 99.14 99.77 97.75 99.94 92.54 96.45 99.66 99.43 99.95 99.92 99.98 99.56
TABLE VII: Ablation study for investigating effect of each part of input embeddings in the scoring network of ${}_{+}$. The symbols n, g, and r represent the node embedding, the graph embedding, and the final relation embedding, respectively.
@@ 4.5 Ablation Studies
TABLE VIII: Ablation study for investigating the effect of each part of different relation correction patterns in ${}_{+}$. The symbol “✓” denotes the types of relation correlation patterns that are taken into account by the RCN module.
@@ 4.5.1 Ablation on relation aggregation and coefficients
In our proposed method, we aggregate the relation embedding $\textbf{r}_{t}$ and neighborhood embedding $\textbf{r}_{t}^{N}$ to get the final relation embedding $\textbf{r}_{t}^{F}$. We omit the aggregation of neighborhood embedding, that is, we let the output of the relational correlation module be $\textbf{r}_{t}$. We reason on the CCN+ subgraph as an example and call this method “${}_{+}$ w/o RA” for short.
@@
where $\mathcal{N}(r_{t})$ represents the set of neighborhood relations of $r_{t}$. We reason on the CCN+ subgraph as an example and call this baseline “${}_{+}$ w/o RC” for short.
Table VI shows the results on three benchmark datasets. The experiments demonstrate the effectiveness of modeling topology-aware correlations between relations in ${}_{+}$. As correlations between relations are common in knowledge graphs, the relation aggregation in “${}_{+}$ w/o RC” can take advantage of neighborhood relations, which is helpful for inductive link prediction. Our proposed method further distinguishes the correlation patterns and correlation coefficients between relations, which makes the learned embeddings of relations more expressive for inductive link prediction. As we can see, ${}_{+}$ significantly and consistently outperforms “${}_{+}$ w/o RA” and “${}_{+}$ w/o RC” on all the inductive datasets.
@@
TABLE X: Comparison between the $2$-hop TACO, $2$-hop ${}_{+}$, and $3$-hop enclosing subgrah TACT in the triple classification task.
@@
TABLE XI: The hits@10 of the frequency-based method, TACT, and the proposed ${}_{+}$.
@@
${}_{+}$ 95.12 75.93 60.81
TABLE XII: The results for inductive link prediction of GraIL, TACT, and the proposed ${}_{+}$ on the dataset YAGO3-10.
@@
${}_{+}$ 0.930 0.471 0.184
TABLE XIII: The running time of GraIL, TACT, TACO, ${}_{+}$, and NBFNet on the version 1 of the inductive datasets. We measure these methods on the same device for a fair comparison.
@@ 4.5.2 Ablation on the input embeddings
We conduct ablation experiments to get access to the exact effect of each part of embeddings in the inductive link prediction. Table VII shows the results of scoring a triple based on different combinations of embeddings. We can see that any part of the embeddings serves its own distinct effect on the final performance results. When performing inductive relation prediction, using the final relation embedding solely—which is exactly the method of our proposed baseline model ${}_{+}$-base—can get a fairly good performance. This shows that modeling the semantic correlation between relations is beneficial to make the correct relation prediction in the inductive setting. The various embedding combinations will further promote the performance on different datasets. We can set the used embedding combination properly to further improve the performance of TACO to get the best results on different benchmark datasets for inductive link prediction.
@@ 4.5.3 Ablation on Different Relation Correlations
In the proposed method, we take into account all different relation correlations of ${}_{+}$. We conduct ablation experiments to figure out the influence on different relation correlations on ${}_{+}$. By reducing the number of relation correlations and analyzing the impact of each correlation on inductive link prediction, we aim to investigate the effect of relation correlations in detail. As shown in Table VIII, although some datasets exhibit improved performances when considering fewer relation correlation patterns, taking into account all relation correlation patterns leads to more stable and generally acceptable results. In specific datasets, omitting a particular relation correlation pattern, such as the ”H-H” pattern on the FB15K-237-v1 dataset, results in significant performance degradation. Therefore, including all relation correlation patterns in the analysis yields more stable and robust results.
@@ 4.6.3 The Frequency-based Method
We conduct an experiment by ranking relations according to their frequencies and compare ${}_{+}$ with the frequency-based method on the datasets. For the frequency-based method, the returned rank list for every prediction is the same, which is the rank according to the relation frequencies from high to low in the knowledge graph. In other words, the frequency-based method represents a type of data bias in the datasets. As illustrated in Table XI, ${}_{+}$ and TACT significantly outperform the frequency-based method by a significant margin on the datasets. The results demonstrate that the effectiveness of ${}_{+}$ is not due to the data bias of relation frequencies in benchmark datasets.
@@ 4.6.4 Results on YAGO3-10
To demonstrate the effectiveness of our proposed method on a larger knowledge graph with few relations. We conduct experiments on YAGO3-10, which is a subset of YAGO3 [46] and contains 37 relations and 123,182 entities. Table XII shows the results for inductive link prediction of GraIL, TACT, and ${}_{+}$ on YAGO3-10. As we can see, ${}_{+}$ outperforms GraIL and TACT by all the metrics, which demonstrate our proposed method can effectively deal with a larger knowledge graph with few relations and Complete Common Neighbor induced subgraph also helps to improve TACO by preserving more relevant rules within the extracted subgraphs.

### v3

@@ 3.5.3 The CCN subgraph method
As shown in Figure 5(b), compared with the $2$-hop enclosing subgraph, after adopting the $2$-hop CCN subgraph extraction method, isolated common neighbor $e_{2}$ will be preserved as an indicator for the existence of the rule from head entity $u$ to the tail entity $v$ through $e_{2}$. It preserves complete common neighbors and more relevant relations from the subgraph efficiently. And compared with the $3$-hop enclosing subgraph, as shown in Figure 6(b), the $2$-hop CCN subgraph effectively eliminates the irrelevant nodes $e_{7}$, $e_{8}$, $e_{9}$ and their corresponding relations. We summarize the procedure of CCN subgraph algorithm in Algorithm 1. And we call TACO that reasons on CCN subgraph as TACOCCN, abbreviated as TACO.
@@ 3.5.4 The CCN+ subgraph method
The CCN+ method can effectively distinguish the relevant and irrelevant nodes. As Figures 5(c) and 6(c) illustrate, The CNN+ subgraph can properly preserve the relevant rules and eliminate the irrelevant rules. We summarize the procedure of the CCN+ subgraph algorithm in Algorithm 2. And we call TACO that reasons on CCN+ subgraph as TACOCCN+, abbreviated as TACO+.
@@ 4.4.2 Ranking metric
TABLE VI: Ablation results of triples classification on the inductive benchmark datasets extracted from WN18RR, FB15k-237, and NELL-995. “TACO+ w/o RA” represent the baseline that omits the relation aggregation in TACO+. “TACO+ w/o RC” represent the baseline that performs relation aggregation without modeling correlations between relations in TACO+.
@@
TACO+ w/o RA 97.75 97.33 91.95 97.64 92.06 95.83 97.05 97.42 98.06 97.33 96.46 98.26
TACO+ w/o RC 98.82 93.17 90.29 97.85 92.39 95.97 96.69 99.22 98.73 97.32 98.81 98.39
TACO+ 99.14 99.77 97.75 99.94 92.54 96.45 99.66 99.43 99.95 99.92 99.98 99.56
TABLE VII: Ablation study for investigating effect of each part of input embeddings in the scoring network of TACO+. The symbols n, g, and r represent the node embedding, the graph embedding, and the final relation embedding, respectively.
@@ 4.5 Ablation Studies
TABLE VIII: Ablation study for investigating the effect of each part of different relation correction patterns in TACO+. The symbol “✓” denotes the types of relation correlation patterns that are taken into account by the RCN module.
@@ 4.5.1 Ablation on relation aggregation and coefficients
In our proposed method, we aggregate the relation embedding $\textbf{r}_{t}$ and neighborhood embedding $\textbf{r}_{t}^{N}$ to get the final relation embedding $\textbf{r}_{t}^{F}$. We omit the aggregation of neighborhood embedding, that is, we let the output of the relational correlation module be $\textbf{r}_{t}$. We reason on the CCN+ subgraph as an example and call this method “TACO+ w/o RA” for short.
@@
where $\mathcal{N}(r_{t})$ represents the set of neighborhood relations of $r_{t}$. We reason on the CCN+ subgraph as an example and call this baseline “TACO+ w/o RC” for short.
Table VI shows the results on three benchmark datasets. The experiments demonstrate the effectiveness of modeling topology-aware correlations between relations in TACO+. As correlations between relations are common in knowledge graphs, the relation aggregation in “TACO+ w/o RC” can take advantage of neighborhood relations, which is helpful for inductive link prediction. Our proposed method further distinguishes the correlation patterns and correlation coefficients between relations, which makes the learned embeddings of relations more expressive for inductive link prediction. As we can see, TACO+ significantly and consistently outperforms “TACO+ w/o RA” and “TACO+ w/o RC” on all the inductive datasets.
@@
TABLE X: Comparison between the $2$-hop TACO, $2$-hop TACO+, and $3$-hop enclosing subgrah TACT in the triple classification task.
@@
TABLE XI: The hits@10 of the frequency-based method, TACT, and the proposed TACO+.
@@
TACO+ 95.12 75.93 60.81
TABLE XII: The results for inductive link prediction of GraIL, TACT, and the proposed TACO+ on the dataset YAGO3-10.
@@
TACO+ 0.930 0.471 0.184
TABLE XIII: The running time of GraIL, TACT, TACO, TACO+, and NBFNet on the version 1 of the inductive datasets. We measure these methods on the same device for a fair comparison.
@@ 4.5.2 Ablation on the input embeddings
We conduct ablation experiments to get access to the exact effect of each part of embeddings in the inductive link prediction. Table VII shows the results of scoring a triple based on different combinations of embeddings. We can see that any part of the embeddings serves its own distinct effect on the final performance results. When performing inductive relation prediction, using the final relation embedding solely—which is exactly the method of our proposed baseline model TACO+-base—can get a fairly good performance. This shows that modeling the semantic correlation between relations is beneficial to make the correct relation prediction in the inductive setting. The various embedding combinations will further promote the performance on different datasets. We can set the used embedding combination properly to further improve the performance of TACO to get the best results on different benchmark datasets for inductive link prediction.
@@ 4.5.3 Ablation on Different Relation Correlations
In the proposed method, we take into account all different relation correlations of TACO+. We conduct ablation experiments to figure out the influence on different relation correlations on TACO+. By reducing the number of relation correlations and analyzing the impact of each correlation on inductive link prediction, we aim to investigate the effect of relation correlations in detail. As shown in Table VIII, although some datasets exhibit improved performances when considering fewer relation correlation patterns, taking into account all relation correlation patterns leads to more stable and generally acceptable results. In specific datasets, omitting a particular relation correlation pattern, such as the ”H-H” pattern on the FB15K-237-v1 dataset, results in significant performance degradation. Therefore, including all relation correlation patterns in the analysis yields more stable and robust results.
@@ 4.6.3 The Frequency-based Method
We conduct an experiment by ranking relations according to their frequencies and compare TACO+ with the frequency-based method on the datasets. For the frequency-based method, the returned rank list for every prediction is the same, which is the rank according to the relation frequencies from high to low in the knowledge graph. In other words, the frequency-based method represents a type of data bias in the datasets. As illustrated in Table XI, TACO+ and TACT significantly outperform the frequency-based method by a significant margin on the datasets. The results demonstrate that the effectiveness of TACO+ is not due to the data bias of relation frequencies in benchmark datasets.
@@ 4.6.4 Results on YAGO3-10
To demonstrate the effectiveness of our proposed method on a larger knowledge graph with few relations. We conduct experiments on YAGO3-10, which is a subset of YAGO3 [46] and contains 37 relations and 123,182 entities. Table XII shows the results for inductive link prediction of GraIL, TACT, and TACO+ on YAGO3-10. As we can see, TACO+ outperforms GraIL and TACT by all the metrics, which demonstrate our proposed method can effectively deal with a larger knowledge graph with few relations and Complete Common Neighbor induced subgraph also helps to improve TACO by preserving more relevant rules within the extracted subgraphs.